### v1

# Gentle Object Retraction in Dense Clutter Using Multimodal Force Sensing and Imitation Learning

Abstract
Dense collections of movable objects are common in everyday spaces—from cabinets in a home to shelves in a warehouse. Safely retracting objects from such collections is difficult for robots, yet people do it easily, using non-prehensile tactile sensing on the sides and backs of their hands and arms. We investigate the role of such sensing for training robots to gently reach into constrained clutter and extract objects. The available sensing modalities are (1) “eye-in-hand” vision, (2) proprioception, (3) non-prehensile triaxial tactile sensing, (4) contact wrenches estimated from joint torques, and (5) a measure of successful object acquisition obtained by monitoring the vacuum line of a suction cup. We use imitation learning to train policies from a set of demonstrations on randomly generated scenes, then conduct an ablation study of wrench and tactile information. We evaluate each policy’s performance across 40 unseen environment configurations. Policies employing any force sensing show fewer excessive force failures, an increased overall success rate, and faster completion times. The best performance is achieved using both tactile and wrench information, producing an 80% improvement above the baseline without force information.
I Introduction
Cluttered and constrained environments, like densely packed shelves and cabinets, appear commonly in household and commercial settings. People interact with them daily to place, organize, and extract objects without conscious thought using a combination of tactile sensing and limited vision [1, 2]. In such scenarios, there are often no collision-free paths to reach and extract desired objects. Hence, if robots are to operate in these scenarios, they need to embrace contacts, rather than avoid them, as in traditional motion planning [3]. At the same time, robots need to be mindful of interaction forces to avoid damage.
Figure 1: a) An end-effector equipped with soft, triaxial tactile sensors, a suction cup, and a camera. b) The “eye-in-hand” camera view provided as observations to a robot reaching in dense clutter to acquire a target. c) Overhead view (not available to robot) of the scene in (b). Robot’s initial pose (blue arrow) causes objects to jam against the left wall, producing a large contact force (red arrow). Action sequence moves the robot to a new pose (green arrow) that is out of contact while still approaching the target.
Gentle object retraction amidst dense collections of movable objects remains a difficult challenge for most robotic systems, especially in constrained lateral access environments where there is an abundance of visual occlusions, objects cannot be freely pushed aside, and there is no overhead access—as exemplified in Fig. 1c. We hypothesize that multimodal vision and force sensing are essential to enable robot policies that embrace contact when it is necessary, but do not cause damage or unecessarily disrupt the environment.
In this work, we consider the task of gently navigating among dense clutter in a cabinet to reach and extract a target object. To complete the task, the robot has access to: (1) “eye-in-hand” monocular vision, (2) its own proprioception, (3) contact wrenches estimated from joint torques, (4) signals from soft, triaxial tactile array sensors, and (5) a measure of object acquisition obtained by monitoring the vacuum line of a suction cup. We opt to use imitation learning to provide an unbiased evaluation of how useful the force sensing modalities are for this task.
Contributions
We provide an investigation of the relative importance of force-torque and tactile sensing for training imitation learning policies to perform gentle, contact-rich manipulation in constrained clutter. We demonstrate that wrench and tactile information provide non-overlapping performance improvements to robots while performing these tasks. To the authors’ knowledge, this work also enables object extraction from more densely cluttered lateral access scenes than prior work (up to 5252% area occupied).
II Related Work
Investigating the importance of force sensing in multimodal systems for contact-rich tasks is an active area of research. Efforts include representation learning of wrench or tactile information [4, 5], using kinesthetic demonstrations[6, 7, 8] or hand-held, robot free demonstrations[9, 10] for tasks that are hard to teleoperate, and learning to perform bimanual manipulation with visuotactile sensors[11, 12]. These approaches provide investigations of the utility of force sensing for performing dexterous, contact-rich manipulation tasks, but do not explore specifically the problem of highly cluttered environments with contact on non-prehensile surfaces.
Other works perform decluttering or packing tasks on tabletop clutter[13, 14, 15, 16, 17], with some approaches using force sensing as part of their multimodal systems[18, 19]. However, in these works the robot can view the manipulation scenario from the top (similar to the overhead view in Fig. 1c), which does not reflect the visual occlusion encountered in lateral access scenarios. Furthermore, an open tabletop environment avoids the sporadic object jamming that frequently occurs in lateral access scenes where the objects are in an enclosed space.
Several recent efforts enable object retraction from cluttered lateral access scenes by performing mechanical search using a slender end-effector with a suction cup at the end [20, 21, 22, 23], yet these approaches are limited to relatively low clutter density, in part due to the lack of force sensing. Specifically, they do not use non-prehensile tactile sensors, despite the majority of manipulation maneuvers being non-prehensile.
Non-prehensile tactile sensors have been shown to enable object motion category classification in constrained clutter[24], but it is non-trivial to incorporate these motion categories into a motion planning system. Prior work has also enabled blind reaching into cluttered environments by leveraging non-prehensile tactile sensing to reach toward target locations amidst fixed obstacles while minimizing contact forces[25, 26, 27, 28]. This approach, however, limits scene variability and avoids the issue of sporadic changes to object impedance when objects jam in constrained yet movable clutter. One recent work uses non-prehensile tactile information to reach toward target locations in constrained, lateral access scenarios by employing motion primitives to react to contacts[29]. All of these approaches assume that target locations are known a priori and do not support target object extraction. Blind object recognition and retraction have been demonstrated using non-prehensile tactile sensing[30], but this is performed in unconstrained tabletop clutter and does not investigate the interplay of vision and force sensing.
III Methods
To develop an autonomous system that can retract objects from constrained clutter, we develop a custom sensorized end-effector and build on the diffusion policy framework presented in [31], leveraging multiple input sensor modalities.
III-A Hardware
We use a Flexiv Rizon 4 arm with a 3D-printed end-effector for manipulation within the cabinet. The arm is controlled via a Robot Operating System (ROS) architecture that coordinates between the Flexiv software and each of the sensing modalities. The slender end-effector, as seen in Fig. 1a, is equipped with a suction cup, an “eye-in-hand” fisheye camera that provides images as seen in Fig. 1b, and two 49-element triaxial tactile sensors as presented in [32]. The minimum signal change that the tactile sensors can distinguish from noise while navigating the environment is $\approx 0.5\,N$, and the measured triaxial contact forces can be represented with circles and arrows proportional to normal and shear force, respectively, as seen in Fig. 2b, or as an RGB image, as shown in Fig. 2c.
Figure 2: a) Images of the left and right side of the end-effector, along with tactile array coordinate frames, indicating how tactile visualizations map to locations on the end-effector. b) Distributed triaxial force information represented with normal force proportional to circle diameter and shear forces proportional to arrow magnitude and direction. c) Corresponding tactile image with x, y, and z forces mapped to B, G, and R channels, respectively.
Figure 3: Our force-informed diffusion policy network processes both input images—“eye-in-hand” camera and tactile—through its own pre-trained ResNet-18 encoder. The output features are directly concatenated with normalized low-dimensional sensor modalities: wrench, proprioception, and pressure. This forms one observation, which is combined with previous observations and passed into a diffusion head to perform action prediction.
In addition to camera and tactile information, we provide proprioception readings using the robot’s joint encoders to form a quaternion representation of the robot’s tool coordinate pose, which is located at the suction cup. We also use built-in Flexiv commands that leverage joint-torque sensing to determine the dynamics-compensated wrench, represented in the tool coordinate frame. Considering the force components of this wrench (moments are small at the suction cup), the minimum signal change that can reliably be detected above noise due to motion of the arm is $\approx 3.3\,N$.
Inspired by [33], we also provide a grasp acquisition metric derived from pressure readings on the suction line. Specifically, we represent pressure as a binary grasp acquisition value that goes high if the pressure reading drops below 6.9-6.9 kPa (0.07-0.07 bar) gauge pressure.
III-B Training Setup
Using the hardware described in the previous section, we collect raw sensor data from demonstrations and package them into observations for training imitation learning policies using the diffusion policy architecture[31].
For our training pipeline, we choose to use visually embedded tactile information based on previous findings that reasoning directly on raw high-dimensional tactile information yields worse performance than reasoning on visually embedded tactile information [5].
Thus, we represent tactile information as a $20\times 5$ pixel image, with each pixel representing one taxel on the end-effector and the three axes of force mapped to the three color channels of the image. First, each tactile array is padded with a single zero-force taxel to make an even 50 taxels. Then, the x forces are mapped from a range of [1N,1N][-1\,N,1\,N] to pixel intensities $[0,255]$ on the blue channel. Similarly, the y forces on the left are mapped from a range of [1N,1N][-1\,N,1\,N] to $[0,255]$ on the green channel. The y force range for the right sensor is flipped so that upward on the end-effector corresponds to increasing values in both cases. As a result, 0 N of shear force corresponds to an intensity of $127$ on both the blue and green channels. The normal z forces are then mapped from [0N,5N][0\,N,-5\,N] to $[0,255]$ on the red channel. Fig. 2 shows an example of how triaxial force information gets mapped to tactile images.
Fig. 3 shows the training and inference pipeline used in our system. The $128\times 128$ pixel “eye-in-hand” camera images are fed through a designated ResNet-18 vision encoder, while the $20\times 5$ tactile images are fed through a separate ResNet-18 encoder. ResNet was found to be the most effective pre-trained visual encoder for use on tactile information in prior work [5]. After normalization, pose, pressure, and wrench are concatenated and fed as separate, low-dimensional inputs to the diffusion policy framework. Diffusion policy then outputs actions that can be executed by the robot controller.
IV Data Collection
To build a dataset of gentle object retraction in clutter, we design a system to generate randomized cluttered environments, outline success and failure metrics for our task, and use a teleoperation setup for gathering expert demonstrations.
IV-A Environment Setup
[…]
Figure 4: a) Schematic of randomly generated configuration of cluttered objects with corresponding b) overhead view of the physical configuration, c) external view of the scene, and d) “eye-in-hand” camera view provided to the robot. Each scene has 25-28 obstacles of 4 types—blue, green, black, and yellow—and one of 3 possible red target objects.
IV-B Task Setup
The task involves starting at a home position in front of the cabinet shelf with a goal of gently navigating through the cluttered environment to acquire a target object and bring it out of the shelf. Successful completion of the task requires locating the correct target object (the only red object in the scene), retracting it to the home position, and dropping it into the region outlined in black (visible beneath the suction cup in the foreground of Fig. 4c,d). We require that the task is completed within 120 seconds.
This task fundamentally requires embracing contacts as there is almost never a contact-free path to the target object, yet the constrained nature of the environment means that objects are likely to impede robot motion, which can cause damage to the environment or the robot if the non-prehensile pushing maneuvers are not performed carefully. To ensure safe behaviors, we require that interactions limit the momentum transfer to the environment. In other words, motions cannot surpass an excessive net force impulse, $I_{net}$
$I_{net}=F_{net}\delta t_{react}$ (1)
where $F_{net}$ is the L-2 norm of the first 3 elements of the measured wrench and $\delta t_{react}$ is the time given to both demonstrator and policy to react to the contact. We also require that motions cannot exceed a peak force impulse, $I_{peak}$
$I_{peak}=F_{peak}\delta t_{react}$ (2)
where $F_{peak}$ is the maximum L-2 norm of the triaxial tactile forces across all taxels.
We implement two distinct impulse thresholds because the peak force is the metric most likely to damage an object, but the wrench measurements are unaware of individual contact magnitudes when there are multiple simultaneous contacts. The peak tactile force, however, does not have full coverage of the robot arm and will be unaware of some contacts, thus a net force threshold is also needed.
We determine $F_{net}$ and $F_{peak}$ values for our task by teleoperating the robot into 2 fragile objects—a plastic cup and the yellow cardboard tea box used in our scenes—until substantial damage occurs. In the case of the tea box, operation was stopped once the cardboard buckled or tore and in the case of the plastic cup, operation was stopped once the cup cracked. We repeat this process 3x for each, then extract the largest value of net force and peak tactile force. We then average these values across all trials and round up to the nearest integer, resulting in a net force magnitude threshold, $F_{net}$, of $26$ N and a peak tactile magnitude threshold, $F_{peak}$, of $6$ N. While we recognize that these forces are not applicable to all possible tasks, they serve as reasonable thresholds to determine gentleness as it pertains to our task and environment.
We select $\delta t_{react}$ according to the standard number of action steps, $n_{act}=8$, in the diffusion policy framework, in combination with the sampling rate, $f_{s}=10\,Hz$, used in our system. In this way, $\delta t_{react}=0.8\,s$ enables the autonomous policies to properly react to contacts in time. Qualitatively, the demonstrator also needed this reaction time in order to properly avoid exceeding the impulse thresholds during teleoperation.
Through this process, we establish the excessive impulse task constraints of $I_{net}<20.8\,Ns$ and $I_{peak}<4.8\,Ns$.
IV-C Demonstration Setup
Figure 5: a) Demonstrator using a spacemouse to control robot motion while observing b) visual, tactile, excessive force warnings, and pressure information provided on screen. The contact on the right side of the end-effector, which is unseen in the camera view, is causing a large peak force as indicated by the red rectangle. The demonstrator does not have an external view of the physical scene.
Using a 3Dconnexion SpaceMouse® Compact to teleoperate the robot, we collect demonstrations on randomized cluttered scenes generated through the process described in Section IV-A. As seen in Fig. 5, the demonstrator has access to the “eye-in-hand” camera view, the tactile images as described in Fig. 2, and pressure sensing, as well as an additional “excessive force” image which visually encodes whether the robot is approaching $F_{net}$ and $F_{peak}$. The left and right sides of this excessive force image (shown in Fig. 5b) represent peak tactile force from the left and right sides of the end-effector while the center portion corresponds to the net force magnitude computed using the joint torques. In these portions, we display grayscale values ranging from black to white as the corresponding force values range from 0% to 8080% of $F_{net}$ and $F_{peak}$. Beyond 8080%, we display a red warning in the corresponding portion to enable the demonstrator to respond. This visualization system was empirically tuned to provide the demonstrator with enough warning to avoid excessive force failures while avoiding false positives as necessary non-prehensile maneuvers are performed.
V Experiment
To investigate the hypothesis that robots need force sensing for gentle object retraction in constrained clutter, we conduct a force ablation experiment by training and evaluating four policies with varying access to wrench and tactile information.
V-A Force Ablation Experiment
Using the setup described in Section IV-C, we gather 100 demonstrations, gently navigating the scene to acquire the red target object. If either $I_{net}$ or $I_{peak}$ is exceeded, the demonstration automatically ends and the scene has to be re-collected by the demonstrator.
With these demonstrations and the architecture described in Section III-B, we trained four policies with varied access to wrench and tactile information to investigate their importance for this task: 1) the baseline policy masks out (replaces with zero values) wrench and tactile information, 2) the wrench-informed policy masks out tactile information, 3) the tactile-informed policy masks out wrench information, and 4) the force-informed policy has both wrench and tactile information included. Apart from this masking, all policies are trained using the same architecture and hyper-parameters, with each policy trained for $200$ epochs.
To ensure that the policies have access to the same information they are trained on, we similarly mask out the wrench or tactile information during evaluations for each policy. We first generate a randomized scene as described in Section IV-A, then evaluate the performance of each of the 4 policies on the same scene according to the success and failure criteria described in Section IV-B.
V-B Results
The results of $40$ evaluations on each policy are summarized in Fig. 6. The first item of note is that each policy that has access to wrench or tactile information in addition to vision significantly outperforms the baseline policy. This can be seen in Fig. 6a, which displays the number of successes, time-out failures, and excessive force failures for each policy. The best performance is demonstrated by the force-informed policy, where we see an 8080% relative improvement above the baseline. The improved performance for all policies with access to either force modality appears to be largely due to a substantial reduction to excessive force failures.
Figure 6: Performance results for 40 evaluations of the force ablation experiment. a) Proportions of successes, time out failures, and excessive force failures for each of the policies. * indicates $p<0.05$ in a two-sided pairwise z-test against the baseline. indicates $p<0.05$ in a two-sided pairwise z-test against the force-informed policy. b) Success times for each policy with error bars showing standard error. Each policy with any force modality improved success rate and accomplished the task faster.
The force-informed policy appears to derive its improved performance largely from the ability to avoid time-out failures as seen in Fig. 6a. In contrast, the significantly higher proportion of timeout failures for the remaining policies often occurred due to pulling away from objects in the absence of large contact forces, failing to focus on the target after it has been isolated, and even retracting without the target object. The baseline policy exhibits the highest timeout failure rate.
Finally, as seen in Fig. 6b, all policies informed by wrench or tactile information accomplished the task faster, with the wrench-informed policy being fastest.
Figure 7: Heatmaps of resulting motion after observing net and peak forces for both the baseline and force-informed policies. a) Motion performed by the baseline policy as a function of observed net forces. b) Resulting motion versus observed peak forces for the baseline policy. c) Resulting motion versus observed net forces for the force-informed policy. d) Resulting motion versus observed peak forces for the force-informed policy. The black dashed lines represent the excessive force values used for our task. The baseline policy displays a higher proportion of large force, low motion datapoints than the force-informed policy, indicating a failure to react to contacts.
To gain a better understanding of how the policies have learned to react to contact forces, we extract the observed forces and resulting displacements of the end-effector for all timesteps of all evaluations of the baseline and force-informed policies. We compute the resulting displacement over a subsequent duration of 0.8s0.8\,s to allow for 1 full inference cycle to enable proper reactions.
The observed motion versus net forces and peak tactile forces can be seen in Fig. 7, where cells toward the bottom right of each plot correspond to a large observed force but little resulting motion—indicating a failure to respond. For reference, the values of $F_{net}$ and $F_{peak}$ are plotted as vertical lines. Ideally, there should be no cells that correspond to little or no motion when forces exceed or approach these values, since remaining in place would maintain the excessive force and cause an impulse failure. As seen in Fig. 7a,b, the baseline policy contains a substantial number of high force, low motion datapoints, but the force-informed policy in Fig. 7c,d avoids low motion as the observed forces increase.
We also note the number of failed reactions by tallying the total instances where less than 0.5cm0.5\,cm of displacement is seen after observing either a net force that exceeds $F_{net}$ or a peak tactile force that exceeds $F_{peak}$. The baseline policy fails to react to large contact forces 412 times in contrast to the force-informed policy seeing only 22 failed reactions.
VI Discussion
When attempting to accomplish object retraction in constrained clutter using imitation learning, we have demonstrated that the addition of force sensing modalities intuitively improves a policy’s ability to be gentle. Beyond this, we note that each policy in our experiment which lacked any force modality that the demonstrator used to react saw a significant increase in timeout failures. These timeout failures appear to be the result of an inability to correctly identify the cause-effect relationship between observations and expert actions.
Previous work establishes a framework to interpret causal relationships in imitation learning by supposing that there are disentangled factors of the observation data, some of which are true causes for expert actions and some of which are nuisance factors that coincidentally correlate with expert actions[34]. This prior work demonstrates that the addition of information—specifically looking at information about the history of actions—can harm policy performance since there are strong correlations between prior actions and expert decisions, but the true causes are due to current observations (i.e. the additional information in the history contains a significant amount of nuisance factors). In this way, removal of nuisance factors from the observation data can reduce causal confusion in imitation learning.
We suspect that in our experiments, the baseline policy has overfit to nuisance factors in the absence of force information, which was the true cause of demonstrator reactions. For example, consider the instances where a quick motion to the right was performed after the demonstrator observed a large peak tactile force on the left. Further suppose that in many of these instances, a specific face of the yellow objects used in our scenes coincidentally appeared in the top right of the camera images leading up to these reactions—thus, we have a nuisance factor in our observation data. During rollouts, when this visual observation reappears and no longer correlates to a large force on the left, a policy which lacks the true, force-informed causes of such reactions is more susceptible to the nuisance factor and performs an inappropriate reaction to the right. The accumulation of such reactions may be the cause of the increased timeout failures for all policies which lacked any force modality used by the demonstrator. This indicates that causal confusion can be remedied not only by removing observation data that is ridden with nuisance factors but also by adding observation data that contains true causes.
As robots increasingly delve into unstructured environments, we should expect that conventional visuo-proprio policies will suffer due to an inability to react to unexpected contacts not only at the grasping surfaces of the robot, but also along non-prehensile surfaces. This work highlights the need of both multi-modality and whole-body sensing, especially if we seek to enable robust manipulation in unstructured human environments.
VII Conclusion & Future Work
In this work, we demonstrate the ability to use imitation learning to train autonomous policies for gently navigating and retracting objects from extremely variable, densely cluttered lateral access environments. We find that in contrast to a baseline policy without access to wrench and tactile information, any policy with access to one or both force modalities demonstrates an improvement in overall success rate, a reduction in excess force impulse failures, and an improvement to the average success time. The fully force-informed policy shows the best performance, with an improvement of 8080% over the baseline.
Future work in this area includes improving demonstration data by providing the demonstrator with a kinesthetic haptic feedback teleoperation device and a haptic sleeve (e.g. as in [35]) to supply wrench and tactile feedback, respectively. We believe this would enable demonstrators to respond more naturally to contact forces, likely leading to more effective trained policies. The hardware could also be adapted to better capture the potential of non-prehensile tactile sensing by building in more degrees of freedom for reacting to tactile stimuli. The end-effector could be replaced with a multifingered manipulator, though this would require more complex tactile sensing geometries and the development of a new grasp acquisition metric.
Additional improvements to the system could involve adding an external visual stream that captures an isometric view of the scene or using a pre-trained segmentation network to identify arbitrary objects rather than only red targets. The network architecture could be improved by fine tuning policies with reinforcement learning as in [36], incorporating more reactive diffusion policy architectures as in [37], or building a custom tactile encoder from play data as in [5]. The proposed imitation learning framework could also be compared against a system that implements a model-based, low-level force-informed controller as in [38] integrated with learned motion planning using vision and proprioception.
More broadly, this work represents a call to move beyond robot interaction in mere visual clutter and to embrace the degree of physical clutter encountered on shelves and in cabinets in homes and commercial settings. By learning from the hardware, sensing, and reactive strategies that humans use on a daily basis, we can enable robots to perform robust, contact-rich manipulation in unstructured environments.
ACKNOWLEDGMENT
Toyota Research Institute provided funds to support this work. D. Brouwer was supported by a Stanford Graduate Fellowship. The authors thank Rick Cory, Hongkai Dai, Hojung Choi, Marion Lepert, and Claire Chen for insightful discussions.